Suppose the functions f, g, h and l are as follows:

F(x)=4x^2-5x+1

g(x)=2√2-x/2-x

h(x)=-1/2x+3

l(x)= log4(x+3)-log4(x-2)

3.1 write down Df and solve the inequality f(x)<=0

3.2 Write down Dg, Dh and Dg+h.

3.3 Solve the equation g(x)=-4

3.4 Solve the equation 4^h(x)=8

3.5 Write down Dl and solve the equation l(x)= 1/2


1
Expert's answer
2021-06-18T11:48:58-0400
"f(x)=4x^2-5x+1"

"g(x)=\dfrac{2\sqrt{2-x}}{2-x}"

"h(x)=-\dfrac{1}{2x+3}"

"l(x)=\log_4(x+3)-\log_4(x-2)"

3.1


"f(x)=4x^2-5x+1"

"Df:(-\infin, \infin)"


"f(x)\leq 0=>4x^2-5x+1\leq 0"

"4x^2-4x-x+1\leq 0"

"4x(x-1)-(x-1)\leq 0"

"4(x-1)(x-\dfrac{1}{4})\leq 0"

"\dfrac{1}{4}\leq x\leq 1"

"x\in\big[\dfrac{1}{4}, 1\big]"

3.2


"g(x)=\dfrac{2\sqrt{2-x}}{2-x}"


"2-x\geq0\ and \ 2-x\not=0"

Then "x<2"

"Dg: (-\infin, 2)"


"h(x)=-\dfrac{1}{2x+3}"


"2x+3\not=0"

Then "x\not=-\dfrac{3}{2}"


"Dh: (-\infin, -\dfrac{3}{2})\cup(-\dfrac{3}{2}, \infin)"



"(g+h)(x)=\dfrac{2\sqrt{2-x}}{2-x}-\dfrac{1}{2x+3}"

"D(g+h): (-\infin,- \dfrac{3}{2})\cup(-\dfrac{3}{2}, 2)"


3.3


"g(x)=4"

"\dfrac{2\sqrt{2-x}}{2-x}=4, x<2"

"\sqrt{2-x}=\dfrac{1}{2}"

"2-x=(\dfrac{1}{2})^2"

"x=\dfrac{7}{4}"

3.4


"4^{h(x)}=8"

"4^{h(x)}=4^{3\/2}"

"h(x)=\dfrac{3}{2}"

"-\dfrac{1}{2x+3}=\dfrac{3}{2}"

"6x+9=-2"

"6x=-11"

"x=-\dfrac{11}{6}"

"x=-\dfrac{11}{6}"


3.5

"l(x)=\log_4(x+3)-\log_4(x-2)"

"x+3>0\ and\ x-2 >0"

"x>2"

"Dl: (2, \infin)"


"l(x)=\dfrac{1}{2}"

"\log_4(x+3)-\log_4(x-2)=\dfrac{1}{2}, x>2"

"\log_4\dfrac{x+3}{x-2}=\dfrac{1}{2}"

"\dfrac{x+3}{x-2}=4^{1/2}"

"\dfrac{x+3}{x-2}=2"

"x+3=2x-4"

"x=7"

"x=7"